OpenStudy (anonymous):

help me Solve -3x2 + 30x - 90 = 0. answer choices: x equals 5 plus or minus 2i square root of 5 x equals 5 plus or minus i square root of 5 x equals 10 plus or minus 2i square root of 5 x equals 10 plus or minus i square root of 5

OpenStudy (anonymous):

lets move everything to the other side to make it positive so that it would be easier to work with , so it would : 3x^2 - 30x + 90 = 0 we can further simplify this and make our life easier by dividing everything by 3 since these are all multiples of 3 so : \[\frac{ (3x^2 - 30x + 90 = 0 ) }{ 3 } => x^2 - 10x +30 = 0 \]

OpenStudy (anonymous):

do you know how to continue with this or should i explain further?

OpenStudy (anonymous):

can you explain further

OpenStudy (anonymous):

do you understand everything till now?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

@Purplerainbowcherry

OpenStudy (anonymous):

okay, so now we have -10x and 30, what you have to do now is to find two numbers which when multiplied will give you 30, and when added will give you -10

OpenStudy (anonymous):

I have now \[10 \pm \sqrt{-20}\]

OpenStudy (anonymous):

divided by 2 to add on to that

OpenStudy (anonymous):

wait what? how did you divide by 2? there are no more common multiples x^2 - 10x +30 =0 is the simplified expression

OpenStudy (anonymous):

i put it in the quadratic formula

OpenStudy (anonymous):

oohhhhhhh okay I got what you did, but see, there is a simpler way to do it :) use the quadratic formula only when you have big numbers that would take too long to simplify or that cannot be simplified, as it takes more time

OpenStudy (anonymous):

x equals 5 plus or minus i square root of 5 this is the answer I got

OpenStudy (anonymous):